LHC 2012 - latest news

Week 44: Machine coordinators: Mike Lamont, Jan Uythoven

Goal of this week: Luminosity production.

Date

Summary

Link to report from morning meeting

Monday 5 Nov
  • 00:02 Start filling
  • 01:31 Stable Beams fill 3265. Initial lumi 7.0e33 cm-2s-1.
08:30 meeting
Sunday 4 Nov
  • 00:55 Beams lost due to possible spurious BLM trigger point 7.
    Fill 3262 delivered 105 pb-1 in 4h42m
  • 03:59 Stable beams fill 3262, initial lumi 7.0e33 cm-2s-1
  • 04:27 Beams lost due to instability B1. Delivered 11 pb-1 in 0h27m
  • 05:25 Injecting, problems B2 stability, suspect ADT.
  • 05:54 No beam from SPS
  • 07:14 SPS back again, checking ADT at injection.
  • 09:30 Intervention on ADT, set-up of ADT with 6 indivs.
  • 11:30 Ready to inject for physics.
  • 12:54 Stable beams fill 3264, initial lumi 7.1e33 cm-2s-1.
  • 14:25 Beams lost due to spurious BLM trigger L1. 33 pb-1 in 1h31m.
  • 15:27 PS Booster problem.
09:00 meeting 
Saturday 3 Nov
  • 09:04 Beams dumped: Sector 67 trip, RB.A67 ST_ABORT_PIC.
    Integrated lumi 143 pb-1 in 9h25m.
  • 10:00 Access for ALICE and P6 for MPE piquet and QPS, micro switch changed.
  • 13:45 Pre-cycle
  • 14:30 Cryogenics problem IP2
  • 16:10 Cryo OK
  • 16:40 Injecting physics beam
  • 16:42 Successful start-up of daughter number 2
  • 19:15 Stable beams fill 3261, initial lumi 7.3e33 cm-2s-1
09:00 meeting
Friday 2 Nov
  • 02:50 Stable beams Fill 3252 Initial luminosity: 7.12e33 cm-2s-1
  • 06:00 Beams lost to QPS trigger on RCBXV1.R1 Re-armed by OP
  • Fill 3252: 65 pb-1 in 3h13m
  • 07:15 Injecting physics beams
  • 08:45 Stable beams Fill 3253 Initial luminosity ~7.3e33 cm-2s-1
  • 09:45 Increase RF voltage from 10 MV to 12 MV
  • 12:00 ADT to wideband settings
  • 12:38 Beams dumped, lost BLM crate L8. 79 pb-1 in 3h53m
  • BLM card change and BLM CPU change. Some access.
  • 17:48 Start pre-cycle
  • 19:30 Transfer line steering
  • 20:29 Beams dumped on last injection B2, LHCb BCM
  • 22:09 Filling for physics
  • 23:39 Fill 3259. Stable Beams. Initial lumi 6.5e33 cm-2s-1
08:30 meeting
Thursday 1 Nov
  • 09:15 Fill still in, waiting for SPS before dumping
  • Keep fill 3249 until 12:40: 197 pb-1 in 15 hours
  • 15:55 Back in stable beams Fill 3250 Initial luminosity 7.23e33 cm-2s-1
  • 20:10 Beams lost to spurious QPS trigger on RQX.L1 Heaters fired, lost cryo start and maintain, vacuum valves around IP1 closed
  • Fill 3250: 83 pb-1 in 4h17m
  • 22:55 Cryo Maintain back
  • 23:10 precycle triplet
08:30 meeting
Wednesday 31 Oct
  • 0:004 Beams dumped, ACSLine4B1: driver fault
    Integrated lumi 85 pb-1 in 4h27m.
  • 01:30 Filling
  • 02:00 MKI.8C vacuum spike - reset and extended softstart
  • 04:15 Stable beams fill #3244, initial lumi 7.34e33cm-2s-1
  • 06:26 Beams lost due to QPS - RQ4.L1, spurious self trigger of heaters. 50 pb-1 in 2h14m
  • Ramp down
  • 09:00 Access in shadow of fibre optics cable move and SPS EDF intervention
  • 12:30 Optical fibre relocation finished
  • 12:52 Access finished, followed by ramp with loop closed instead of standard pre-cycle
  • 15:00 Injecting probes
  • 16:00 Injection steering, not evident
  • 21:42 Stable beams, fill #3249, initial lumi 7.1e33 cm-2s-1
08:30 meeting
Tuesday 30 Oct
  • 03:32 Beam dumps two correctors tripped in 5. Int. lumi 80.4/1.1/82.6/6.5 pb-1.
  • Problems refilling: RF crowbar (solved, not understood), PC faults IT.R5.
  • 7:00 Access for EPC, problem solved.
  •  08:40 Pre-cycle IT.R5.
  • 09:34 Injecting probe, fill  #3240.
  • 12:40 Stable beams, intial lumi 6.95e33cm-2s-1
  • 14:06  Beams lost due to QPS trigger. 32 pb-1 in 1h28m
  • 16:00 ADT team adjusting high bandwidth settings
  • 18:10 Injecting for physics
  • 19:35 Stable beam fill #3242, initial lumi 7.04e33cm-2s-1
08:30 meeting
Monday 29 Oct
  • 08:48 V/dM scan CMS end of fill #3236.
  • 10:06 V/dM scan finished, start of RF studies, reducing RF voltage.
  • 11:10 RF voltage reduced from 12 MV to 9 MV.
  • 11:51 Opened TCLs to test for future tests.
  • 12:00 TCL tests finished.
  • 12:08 Beams dumped by operator. Integrated lumi 216/0.20/211/25.1 pb-1.
  • 13:10 Access for experiments, EPC and CV.
  • 15:26 QPS trip of matching section R8, related to EPC intervention on RD2 earlier.
  • 18:00 Bad contact found, start pre-cycle.
  • 18:54 Injecting probes.
  • 19:50 Ramp started fill #3237.
  • 19:50 Beam dumped by LBDS at start of ramp, energy tracking MKD B2 generator L.
  • 21:12 Injecting physics beam fill #3238.
  • 22:55 Stable beams, initial lumiís 6.5e33 cm-2s-1.
08:30 meeting